OpenStudy (anonymous):

An electron in the 4d6 orbital would have: more energy than an electron at the 4f6 orbital more energy than an electron at the 4p6 orbital more energy than an electron at the 5p6 orbital more energy than an electron at the 5f6 orbital

OpenStudy (chmvijay):

n+l rule can help you higher the n+l for the given orbital higher its energy . for example 3d and 4s for 3d n=3 and l=2 n=l =3+2=5 similarly for 4s n=4 and l=o n=l=4+0=4 so 3d has higher energy compared to 4s similarly can u find out for your things